OpenStudy (bananas1234):

If p varies directly with q, and p = 10 when q = 5, what is the value of p when q = 20? (Points : 5) p = 1 p = 28 p = 40 p = 100

OpenStudy (shaik0124):

when q =5 p=2q=2*5=10 similarly when q=20 p=2q=2*20=40

OpenStudy (bananas1234):

i'm still kinda confused...

OpenStudy (aripotta):

set up a proportion  cross-multiply and solve for p